OpenStudy (anonymous):

What is the potential difference across 300ohm resistor when a current of 2.10 A flows through it?

OpenStudy (anonymous):

R = V/I 300 = V/2.1 V= 630v

OpenStudy (anonymous):

Do you not need to use 300* e-3 (2.1) ? It still states the problem is incorrect.

OpenStudy (anonymous):

where u get this 300* e-3 (2.1), i don get it.

OpenStudy (anonymous):

That is what it states in my book however when I tried answering it the way you showed me it was still incorrect.

OpenStudy (anonymous):

what is the answer on ur book?

OpenStudy (anonymous):

it didn't give us answer it just shows you how to set the equation which is V=IR

OpenStudy (anonymous):

yeah v=IR so V= 300 * 2.1, it suppose to be the right one i guess

OpenStudy (anonymous):

It's still stating my answer is incorrect

OpenStudy (anonymous):

it didn't give us answer it just shows you how to set the equation which is V=IR

OpenStudy (anonymous):

That is what it states in my book however when I tried answering it the way you showed me it was still incorrect.

OpenStudy (anonymous):

Do you not need to use 300* e-3 (2.1) ? It still states the problem is incorrect.

OpenStudy (anonymous):

post the circuit diagram..

OpenStudy (anonymous):

what is that e stand for?

OpenStudy (anonymous):

@allenry ,if the circuit diagram has been given in your book then post it...

OpenStudy (anonymous):

i guess this kind of question don show any diagram, maybe

OpenStudy (anonymous):

In the problem you wrote here, you list the units on the resistor as being Ohm's. In this case, V = 300 * 2.1 is valid and yields the correct answer. However, you keep insisting that we should use 300 e-3 Ohms, which makes me think the resistor has units of 300 milliOhms.

OpenStudy (anonymous):

@allenri if the resistance is in milliohm then V=300*e^-3*1.2 is absolutely right answer.